If a spring is dropped from a height and while it is in mid air a mass 'm' is kept on it.Is the weight of the mass an external force on the spring

Grade:12

2 Answers

Askiitians Expert Nagaraju
14 years ago
Hi JITIN Hope doing well in your studies. For a freely falling bodies case If our weight of the mass which you put more than the Weight of the spring it will reach the earth quickly compare to the light weight. Please feel free to ask as many questions as you can to get fast answers. Regards Askiitians Experts Nagaraju
Rohith Gandhi
14 years ago

Dear Jitin,

As you know gravitational force on any object is given as , F = mg , so it seems that the heavier mass i.e. the block of mass m that you kept is going to be pulled with strong force than the light spring. But the point is what accelarations do both the masses have, depending on that we can say which one is going faster.

consider spring having mass , m1,

force on it is, F1 = m1Xg => accelaration of the spring downwards = F1/m1 = m1Xg/m1 = g (which doesn't depend on mass of spring)

now consider block of mass m which you kept on the spring,

force on it by the earth = mg, => accelaration of the block downwards = F/m = mg/m = g which is same as the above,

so, both spring and block have same accelaration = g, hence they do not exert any force on each other, hence weight of the mass on spring is zero

this implies that all freely falling bodies have same accelaration which doesn't depend on the mass of the respective bodies.

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best STUDENT NAME !!!




Regards,

Askiitians Experts

Rohith Gandhi, IIT Kgp